FLWO Changelog

This is a running list of all significant changes that we, the staff of FLWO would like to pass along to all users.

Staff can use this link to Add to this changelog!


ADC work complete. Prisms going to correct positions all sky. Updated home pos, replace PA code, swapped A and B prism definitions.
P Berlind, A Colson <tresobs>
Telescope Involved: 1.5m
Date changed: 08/18/20
Submitted - Wednesday, September 16, 2020 at 04:48:29 (MST)
Marc and I replaced TRES incand cal lamp with spare lamp from cabinet. Counts up 50%, 20k to 30k in extracted spectra. Was expecting more counts, will continue investigating and acquire new spare lamp.
P Berlind <tresobs>
Telescope Involved: 1.5m
Date changed: 07/28/2020
Submitted - Tuesday, July 28, 2020 at 12:46:46 (MST)
Set ADC software home position to 202 to match install.
P Berlind <tresobs>
Telescope Involved: 1.5m
Date changed: 07/25/2020
Submitted - Monday, July 27, 2020 at 04:12:43 (MST)
Primary mirrors of both telescopes CO2'd, 48" filters air dusted
Mike
Telescope Involved: 60" & 48"
Date changed: 03/13/2020
Submitted - Friday, March 13, 2020 at 02:36:11 (MST)
Keplercam warmed the night of Sat 02/01. It had a normal LN2 fill at 17:30 on 02/01. The cause remains unexplained. I came on Sun 02/02 and pumped and cooled the dewar. On 02/03 it reached -96C, but it was clear that the CCD had severe measles. I let the CCD warm back up through 02/04. On 02/05 I conducted a bake. I pumped and cooled on 02/06. After it reached -96 images were much better. Ran well the night of 02/06. Marc Lacasse was a great help.
Emilio Falco <Falco>
Telescope Involved: 1.2m
Date changed: 02/07/2020
Submitted - Friday, February 07, 2020 at 13:34:14 (MST)
I replaced the dome encoder drive belt. Used 208 teeth version out of 3x212 and 3x208. Old one broke in my hands as I was pulling it out.
Emilio Falco <falco>
Telescope Involved: 1.2m
Date changed: 01/24/20
Submitted - Friday, January 24, 2020 at 12:43:57 (MST)
Emilio and I cleaned TRES front-end tiptilt mirror.
Perry Berlind <tresobs>
Telescope Involved: 1.5m
Date changed: 01/21/2020
Submitted - Wednesday, January 22, 2020 at 15:39:52 (MST)
Daniel Y had to realign the dome once last night. Damon got us some friction tape. I applied that to the wheel today. Dome went back to home (using home dome in tcs) after several revs CCW and CW.
Emilio Falco <Falco>
Telescope Involved: 1.5m
Date changed: 01/14/2020
Submitted - Tuesday, January 14, 2020 at 12:20:13 (MST)
Also, found dome encoder mounted on plate was held by only 1 screw. I found one and a nut on the floor that may have come off the 2nd screw hole, put it back on. May help!
Emilio Falco <Falco>
Telescope Involved: 1.5m
Date changed: 01/13/2020
Submitted - Monday, January 13, 2020 at 15:06:28 (MST)
Over past month,dome was not tracking well. After several iterations of adding electrical tape to the wheel, today it seems back to good after 2 full revs each CW and CCW. I tightened the bracket, both with the spring and by adding 2 C-clamps on the side of the broken hinge. Reset the counts/deg back to initial 32.094.
Emilio Falco <Falco>
Telescope Involved: 1.5m
Date changed: 01/13/2020
Submitted - Monday, January 13, 2020 at 15:01:51 (MST)
Damaged, now repaired RA motor came back after DFM tested it. Damon Burke installed it, back to normal. Old RA motor and parts for mounting it are in the telescope pedestal.
Emilio Falco <Falco>
Telescope Involved: 1.5m
Date changed: 01/10/2020
Submitted - Monday, January 13, 2020 at 14:57:14 (MST)
On 12/30/19, the dome started blowing fuses. I worked with Tom Gerl. Pulled the DCX302C board (>20 years old, no drop-in replacement). It has 5 SCRs. I found 2 of them had shorts between anode and cathode, a bad sign. Ordered replacements, with spares. Tom removed the old ones and soldered the new ones in today. It worked! Larry helped today, tightened the encoder belt.
Emilio Falco <Falco>
Telescope Involved: 1.2m
Date changed: 01/08/2020
Submitted - Wednesday, January 08, 2020 at 14:33:12 (MST)
TRES had been trying to warm up. Gil added a 3pm fill to the 9am fill. The ion pump would start, after turning on the HV it quickly went into protect mode. I tried to turn on the HV several times, each time with a slightly lower pressure, down to about 4E-5 mB. I raised the protect threshold from 5 to 14 mA. Eventually, it started holding and the pressure was down to 2E-6 mB. The current was just below 1mA. Will not pump the dewar if that holds. Keeping 2 fills for now.
Emilio Falco <Falco>
Telescope Involved: 1.5m
Date changed: 12/03/2019
Submitted - Wednesday, December 04, 2019 at 10:58:10 (MST)
Another failed video card. This time on flwo48. Old card heated to 136C and caused flwo48 to shutdown hard (tut it still was powered). The card fans were not turning when we looked in the case. Replaced with 4GB dual GEForce GT360, last spare, ordering 4 more.
Emilio Falco <Falco>
Telescope Involved: 1.2m
Date changed: 11/22/19
Submitted - Sunday, November 24, 2019 at 15:12:49 (MST)
New RA motor failed, gone to DFM and on to MTI for repair. Turn-around is 1 month. A new motor is $2300, 4-6 weeks. For a temporary solution, Damon Burke of DFM came out and installed old RA motor with encoder. System workable with FAST but still needs adjustments. Also, handpaddle S button stopped working. Reported to DFM.
Emilio Falco <Falco>
Telescope Involved: 1.5m
Date changed: 11/20/19
Submitted - Sunday, November 24, 2019 at 15:08:40 (MST)
With TG direction, changed flwo60 video card, which was causing the system to overload and stop working altogether.
Emilio Falco <falco>
Telescope Involved: 1.5m
Date changed: 11/19/2019
Submitted - Tuesday, November 19, 2019 at 16:03:51 (MST)
Cell chiller stopped cooling. Had Cactus H&C come, Tom Shultz. He replaced a solenoid valve that leaked and recharged with 404 freon, which was all gone.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 11/12/2019
Submitted - Thursday, November 14, 2019 at 11:46:38 (MST)
I found dome wheel 4 had broken off its plate. Larry S welded it and installed it.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 11/05/19
Submitted - Friday, November 08, 2019 at 11:45:40 (MST)
Installed new HA encoder ramp that Ian Huss made. Worked with Perry at the handpaddle. New E HA limit is about 4h.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 11/01/2019
Submitted - Friday, November 01, 2019 at 16:12:33 (MST)
Cleaned the HA drive arc again. Found crud on it, possibly due to moths. Several flew out when I opened the back of the drive arc. Went back & forth several times until no crud appeared. No slips on the night of 10/15.
Emilio Falco <Falco>
Telescope Involved: 1.2m
Date changed: 10/15/2019
Submitted - Wednesday, October 16, 2019 at 12:01:16 (MST)
I cleaned off and re-lubed the wormgear using LubriPlate 3000, per DFM instructions.
Emilio Falco <Falco>
Telescope Involved: 1.5m
Date changed: 09/16/2019
Submitted - Monday, September 16, 2019 at 17:23:36 (MST)
Telescope was seen slipping on 09/03, as back in June 2019. I tightened the drive screws in the front but it seemed a cleaning of the RA drive arc was in order. Although it seems too soon. I tried a few times to remove the cover as we did in June, no dice even with muscle help. Ended up opening the back of the cover on the S end. That gives about 4cm clearance to clean the arc. It was not terrible but it did have a few squashed moths at the ends. It's now clean and on 09/12/2019 there was no slippage.
Emilio Falco <Falco>
Telescope Involved: 1.2m
Date changed: 09/11/2019
Submitted - Thursday, September 12, 2019 at 15:45:44 (MST)
EF Examined Kep shutter controller, replaced a 2A/250V fuse which seemed OK anyway, saw nothing awry. EF then swapped in a 2nd, older controller that works, so Kep is back in action.
Emilio Falco <Falco>
Telescope Involved: 1.2m
Date changed: 09/03/2019
Submitted - Tuesday, September 03, 2019 at 16:36:37 (MST)
EF and WAP cooled and pumped Keplercam on 08/26-28. On 08/28 we found the chip had measles, requiring baking. Baked 08/29-30 (recipe by WAP on the phone). When it was time to take an exposure, found the shutter control box had failed (no lights, no action). Will try to troubleshoot asap.
Emilio Falco <Falco>
Telescope Involved: 1.2m
Date changed: 08/30/2019
Submitted - Sunday, September 01, 2019 at 13:13:45 (MST)
TRES guider failed on attempt to run on the sky on 08/26. GE replaced guider fiber, but it was the wrong one. TG bought 2 fiber pairs of right type. Fiber worked, but guider did not. Finally on 08/30, TG pulled guider card out of PC, blew on it, and at that point it all worked. EF and TG ran the fiber properly. Great save by TG!
Emilio Falco <Falco>
Telescope Involved: 1.5m
Date changed: 08/30/2019
Submitted - Sunday, September 01, 2019 at 13:05:23 (MST)
DFM updated the control system to Galil, installed encoders on HA, DEC and focus. New machine runs Win 7, with TheSky X added. Replaced dome motors and limit switch. Many loose ends left to fix after DFM left.
Emilio Falco <Falco>
Telescope Involved: 1.5m
Date changed: 08/08/2019
Submitted - Sunday, September 01, 2019 at 13:01:47 (MST)
TG and EF opened the side of FAST. It was quite dusty inside. We blew off and vacuumed as much as we could.
Emilio Falco <Falco>
Telescope Involved: 1.5m
Date changed: 16 Jul 2019
Submitted - Wednesday, July 17, 2019 at 15:28:48 (MST)
After many steps to try to eliminate intermittent tracking problems, we finally found the solution with guidance from DFM. WAP, EF and TG removed the cover on the S side to access the RA drive arc. It was very cruddy. We cleaned it off with cleaning solution for optics and rags. The arc looked like new. It will get cruddy again with time. But we started with no tracking problems overnight.
Emilio Falco <falco>
Telescope Involved: 1.2m
Date changed: 06/20/2019
Submitted - Friday, June 21, 2019 at 14:06:42 (MST)
Swapped the RA and DEC position encoders.
WAP <waynep>
Telescope Involved: 48"
Date changed: 06/14/2019
Submitted - Friday, June 14, 2019 at 15:44:51 (MST)
EF, MC, WAP cleaned the ADC using Gabor Furesz's prescription. It had a significant amount of dirt on it that was completely removed. Re-installing the stovepipe was a bear. WAP ended up cutting a new notch to avoid hanging up on the small gear, and that allowed us to re-install. Not sure why this was not a problem before.
Emilio Falco <falco>
Telescope Involved: 1.5m
Date changed: 5/8/2019
Submitted - Thursday, May 09, 2019 at 10:02:04 (MST)
I replaced the Starna He/Ne/Ar in the FAST spectrograph with the spare lamp.
wap <waynep>
Telescope Involved: 60"
Date changed: 05/07
Submitted - Tuesday, May 07, 2019 at 13:13:37 (MST)
Filters inspected and/or cleaned.
Mike
Telescope Involved: 48"
Date changed: 03/15/2019
Submitted - Friday, March 15, 2019 at 07:06:44 (MST)
Primary mirrors CO2'd
Mike
Telescope Involved: 48", 60"
Date changed: 03/15/2019
Submitted - Friday, March 15, 2019 at 07:06:05 (MST)
WAP and EF replaced broken fiber for TRES guider. Working now. The connector to the guider was damaged on 01/10/19 and it broke for good during the changeover on 02/05/19.
Emilio Falco <falco>
Telescope Involved: 1.5m
Date changed: 02/11/2019
Submitted - Tuesday, February 12, 2019 at 16:54:05 (MST)
Replaced Set and Slew buttons on observer's hand-paddle.
W Peters <waynep>
Telescope Involved: 60"
Date changed: 11/23/18
Submitted - Friday, November 23, 2018 at 12:41:26 (MST)
WAP replaced dome motor+speed reducer with a little help from EEF. Took longer than hoped. It was difficult to find replacements and there were several snafus from Arizona Electric Motor that delayed the installation. Having a spare made up.
Emilio Falco <falco>
Telescope Involved: 1.2m
Date changed: 11/19/2018
Submitted - Tuesday, November 20, 2018 at 09:25:22 (MST)
CO2'd primary mirror.
Mike
Telescope Involved: 60"
Date changed: 10/14/2018
Submitted - Tuesday, October 16, 2018 at 03:22:04 (MST)
CO2'd primary mirror
Mike
Telescope Involved: 48"
Date changed: 10/14/2018
Submitted - Tuesday, October 16, 2018 at 03:21:08 (MST)
Over the month, removed the failed ion pump and controller from the TRES spectrograph bench and rack. Replaced both with new replacements. The dewar is holding at 2x10^-6 Torr. The readout looks reasonable but the temperature is only down in the -67 degC region.
Wayne Peters <waynep>
Telescope Involved: 60"
Date changed: 2018 August, 8/27/18
Submitted - Monday, August 27, 2018 at 15:10:49 (MST)
Helped, mostly watched, Mike CO2 the primary. Significant dust came off.
Emilio Falco
Telescope Involved: 2MASS
Date changed: 7/25/2018
Submitted - Thursday, July 26, 2018 at 17:14:55 (MST)
CO2'2 primary mirror.
Mike
Telescope Involved: 60"
Date changed: 06/15/2018
Submitted - Friday, June 15, 2018 at 04:56:38 (MST)
Cleaned filters.
Mike
Telescope Involved: 48"
Date changed: 06/15/2018
Submitted - Friday, June 15, 2018 at 04:56:01 (MST)
CO2'd primary mirror.
Mike
Telescope Involved: 48"
Date changed: 06/15/2018
Submitted - Friday, June 15, 2018 at 04:55:29 (MST)
Replaced "North" switch in the control room hand paddle.
WAP <waynep>
Telescope Involved: 60"
Date changed: 05/12/18
Submitted - Saturday, May 12, 2018 at 15:09:25 (MST)
TG and EF removed stovepipe and cleaned corrector. First, blow then dab with blue. Greatly improved to the eye. Remember ball for stovepipe no longer goes in the old hole!
Emilio Falco <falco>
Telescope Involved: 1.5m
Date changed: 5/9/18
Submitted - Thursday, May 10, 2018 at 14:50:49 (MST)
CO2'd primary mirror.
Mike
Telescope Involved: 48"
Date changed: 05/09/2018
Submitted - Wednesday, May 09, 2018 at 05:19:47 (MST)
CO2'd primary mirror.
Mike
Telescope Involved: 60"
Date changed: 05/09/2018
Submitted - Wednesday, May 09, 2018 at 05:18:51 (MST)
Repaired TRES projector arm failure. After much testing... , I found that the green wire on digital i/o 4 CWW of the MicroLynx was not inserted completely. It lifted out of the connector with no resistance. After reinserting it properly, the projector arm is once again working. Hopefully, this explains the mysterious projector arm hangs of the past.
W Peters <waynep>
Telescope Involved: 60"
Date changed: 5/04/18
Submitted - Friday, May 04, 2018 at 15:33:16 (MST)
On 2/8 once it was mounted, FAST CCD would not read out other than biases. TG and EF with help from Steve Amato traced the problem to no OR and OD voltage in the controller. With Steve's guidance, on 2/9 TG found no voltage where 32V was expected, then a bad wire on the connector to the CCD power supply. He re-soldered it to its pin and got 32V. Then, the CCD started working normally again. Thanks, all!
Emilio Falco <falco>
Telescope Involved: 1.5m
Date changed: 2/9/2018
Submitted - Friday, February 09, 2018 at 12:37:14 (MST)
TG with help from support removed Keplercam+topbox and replaced the Keplercam shutter. It's a Prontor E/100, same model as the failed one, but slightly different. The shroud needs to be modified to fit it, but time was too tight to do it, so we are w/o a shroud for now. Does not have a significant impact, but we'll adjust it the next chance we have. The replacement shutter has not failed after several nights of operation. We opened the failing shutter, cleaned it. It would not fail at room temp, but did once we cooled it with CO2 snow. There is a raised, shiny section of the internal ring that the leaves ride on. That may be the problem. It is going to Cambridge for a refurb. Schneider does not make the Prontor anymore and none can be found online.
Emilio Falco <falco>
Telescope Involved: 1.2m
Date changed: 01/24/2018
Submitted - Sunday, January 28, 2018 at 13:36:51 (MST)
CO2'd primary mirror.
Mike
Telescope Involved: 60"
Date changed: 01/26/2018
Submitted - Saturday, January 27, 2018 at 06:52:45 (MST)
CO2'd primary mirror
Mike
Telescope Involved: 48"
Date changed: 01/26/2018
Submitted - Saturday, January 27, 2018 at 06:52:25 (MST)
TG came up early to look at the Keplercam shutter, which failed the previous night, producing biases for all exposures. It was intermittent, as it started working after several tries in the AM today. During troubleshooting, the only thing TG found is a wire on the shutter controller that was loose and came off easily. He re-soldered it and the shutter has been operating normally through a couple of sequences of dome flats. We'll call it good for now, but there may be some other problem.
Emilio Falco <falco>
Telescope Involved: 1.2m
Date changed: 01/19/2018
Submitted - Friday, January 19, 2018 at 13:33:31 (MST)
TG+EF switched to spare (older) flat lamp power supply. Dome flat levels rose from <10K to >13K. Unclear why newer supply had drifted down, but we'll stay with the older one for now.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 1/17/2018
Submitted - Wednesday, January 17, 2018 at 14:25:06 (MST)
Mirror CO2'd, filters air-dusted, electrical tape re-applied to gap between ccd dewar and topbox.
Mike
Telescope Involved: 48"
Date changed: 10/27/2017
Submitted - Friday, October 27, 2017 at 15:23:35 (MST)
Mirror CO2'd
Mike
Telescope Involved: 60'
Date changed: 19/26/2017
Submitted - Friday, October 27, 2017 at 15:22:23 (MST)
Replaced the audio taper potentiometer with a linear taper potentiometer in the FAST guide camera gain controller.
Wayne Peters <waynep>
Telescope Involved: 60"
Date changed: 10/19
Submitted - Thursday, October 19, 2017 at 16:08:19 (MST)
Posted new M1 and M2 reflectivity plots, data taken with Minolta on 9/8/17.
Emilio Falco <falco>
Telescope Involved: 1.2m
Date changed: 9/20/2017
Submitted - Wednesday, September 20, 2017 at 16:40:31 (MST)
CO2'd primary mirror.
Mike
Telescope Involved: 60"
Date changed: 09/18/2017
Submitted - Monday, September 18, 2017 at 05:19:16 (MST)
Had Steve Brown (Support) tighten up the telescope chamber double doors/jams to reduce light leaks.
WAP <waynep>
Telescope Involved: 48"
Date changed: 9/14/2017
Submitted - Thursday, September 14, 2017 at 14:23:47 (MST)
Installed the new FAST guide camera gain controller.
WAP <waynep>
Telescope Involved: 60"
Date changed: 9/13/2017
Submitted - Thursday, September 14, 2017 at 14:21:32 (MST)
Wrapped Keplercam with reflective insulation.
Emilio Falco <falco>
Telescope Involved: 1.2m
Date changed: 6/28/2017
Submitted - Wednesday, June 28, 2017 at 15:02:15 (MST)
Mirror CO2'd.
Mike
Telescope Involved: 60"
Date changed: 04/20/2017
Submitted - Tuesday, May 09, 2017 at 21:41:19 (MST)
MIrror CO2'd and filters cleaned.
Mike
Telescope Involved: 48"
Date changed: 04/20/2017
Submitted - Tuesday, May 09, 2017 at 21:40:47 (MST)
WAP and EEF replaced all 4 dome flat lamps, same type as before. Clear improvement, shortened dome flat times.
Emilio Falco <falco>
Telescope Involved: 1.2m
Date changed: 4/14/17
Submitted - Friday, April 14, 2017 at 17:03:09 (MST)
Gil tested the new guider stage that WAP replaced, the night of 3/3. After swapping cables 1 and 3 (x<->y), it worked well. Even better than before, a very good thing. Congrats!
Emilio Falco <falco>
Telescope Involved: 1.5m
Date changed: 3/4/17
Submitted - Saturday, March 04, 2017 at 18:00:18 (MST)
Over the last week I removed the TRES guider tip/tilt stage from the front-end, removed the mirror from the stage, transferred the mirror to the new stage, and re-assembled the front-end. Emilio, and Steve Brown helped at various stages of the work.
W Peters <waynep>
Telescope Involved: 60"
Date changed: 03/02/17
Submitted - Friday, March 03, 2017 at 15:12:12 (MST)
Filters cleaned and/or inspected.
Mike
Telescope Involved: 48"
Date changed: 01/01/2016
Submitted - Monday, January 02, 2017 at 00:08:10 (MST)
CO2'd primary mirror
Mike
Telescope Involved: 48"
Date changed: 01/01/2016
Submitted - Sunday, January 01, 2017 at 23:40:03 (MST)
CO2'd primary mirror
Mike
Telescope Involved: 60"
Date changed: 01/01/2016
Submitted - Sunday, January 01, 2017 at 23:39:37 (MST)
Installed, and successfully tested on the sky, a new Video Scope CCD350-F guide camera, in-house designed and built gain control box, and cabling.
WAP <waynep>
Telescope Involved: 48"
Date changed: 11/22/16
Submitted - Friday, November 25, 2016 at 11:20:18 (MST)
I isolated, and had Larry Simkins plug the light leak in the 48" dome slit.
WAP <waynep>
Telescope Involved: 48"
Date changed: 11/15/16
Submitted - Tuesday, November 15, 2016 at 14:52:45 (MST)
I replaced a faulty breaker in 60" dome rotation control box.
wap <waynep>
Telescope Involved: 60"
Date changed: 11/10/16
Submitted - Thursday, November 10, 2016 at 15:21:09 (MST)
Axial CWs underneath E and S covers adjusted. Moved shafts up and away from stirrups in 6 CWs. Same adjustment on 09/17 for 3 CWs on E side.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 09/19/2016
Submitted - Monday, September 19, 2016 at 15:52:32 (MST)
Adjusted SW radial mirror cell/mirror actuator weight to stop it from contacting the mirror cell.
Wayne Peters <waynep>
Telescope Involved: 48"
Date changed: 09-14-2016
Submitted - Wednesday, September 14, 2016 at 13:39:58 (MST)
Tested collimation. Then tweaked it. First manually, then using aberfit corrections. Looks good. Ensured that guider focus on-axis was good after tel was in focus.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 8/29/2016
Submitted - Tuesday, August 30, 2016 at 11:51:47 (MST)
With PB and MC, adjusted collimation. It looked good after a few turns of the screws. It held N, S, E and W. Pointing was OK.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 8/29/2016
Submitted - Tuesday, August 30, 2016 at 11:49:36 (MST)
Filters cleaned.
Mike
Telescope Involved: 48"
Date changed: 08/04/2016
Submitted - Thursday, August 04, 2016 at 15:14:07 (MST)
CO2'd primary mirror.
Mike
Telescope Involved: 48"
Date changed: 08/04/2016
Submitted - Thursday, August 04, 2016 at 15:13:20 (MST)
Primary mirror CO2'd
Mike
Telescope Involved: 60"
Date changed: 06/03/2016
Submitted - Friday, June 03, 2016 at 06:17:30 (MST)
Primary mirror CO2'd
Mike
Telescope Involved: 48"
Date changed: 06/03/2016
Submitted - Friday, June 03, 2016 at 06:17:03 (MST)
Filters cleaned.
Mike
Telescope Involved: 48"
Date changed: 06/03/2016
Submitted - Friday, June 03, 2016 at 06:16:40 (MST)
On Sat 05/07, FAST had warmed up again. Wayne came up around 9pm to see if he could recover it. He found he would need to bake it and pump it again. He also found 3 screws that were not tight at the bottom of the dewar, where the stinger is inserted. He tightened those. We decided to keep it warm with the pump running until 05/09. Ste kept it at about room temperature overnight. In the morning of 05/09, Wayne started a bake, then the normal recipe to complete the recovery. By 2pm or so, we had good vacuum and the CCD was approaching its operating temperature. When it was cool enough, he checked the alignment, which was very good the first time around, no need to change it. FAST is back on starting tonight.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 05/09/2016
Submitted - Monday, May 09, 2016 at 15:45:35 (MST)
Steve A and Matt S opened the dewar in Cambridge. They found a broken wire between constantan and thermistor. They re-soldered it and closed it up. The dwar is back at the Ridge as of 05/05. Wayne started the process of warming, pumping, and eventually cooling. Plan is to mount FAST back on 05/06.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 05/05/2016
Submitted - Thursday, May 05, 2016 at 19:41:52 (MST)
Could not read resistance (so temp) in FAST dewar. We speculated the thermistor had failed. Sent to Steve A on 05/03, arrived 05/04. SA and Matt Smith found a broken wire between the thermistor and constantan. SA and MS re-soldered it and put the dewar back together. On its way back to FLWO via FedEx 05/04.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 05/04/2016
Submitted - Wednesday, May 04, 2016 at 16:33:55 (MST)
Inspected, blew and used Custom Scientific optics cleaning kit to clean r,z,g,IH. A little better, not great. Filters are degrading.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 04/27/2016
Submitted - Wednesday, April 27, 2016 at 16:07:12 (MST)
Blew i,r,z. Also used first contact on i. Improved!
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 4/26/2016
Submitted - Tuesday, April 26, 2016 at 16:30:23 (MST)
TG+EF lowered dewar, inspected corrector. It was dirty as hell. We blew the dust off. There is some dust on the CCD. After this, dome flats showed MUCH less of the question-mark pattern that's been there for months, but getting worse. We declared victory.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 4/26/2016
Submitted - Tuesday, April 26, 2016 at 14:04:22 (MST)
TG + LS rotated the ADC back to its previous orientation after it was clear it had become worse. No problems as of 4/15.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 4/13/2016
Submitted - Friday, April 15, 2016 at 12:35:29 (MST)
Wayne pointed tel. Larry, Ted, assisted. Removed stovepipe. EF cleaned ADC using 50/50 methanol/acetone per Gabor's instructions online. Looks much better. Turned ADC assy to what may be the nominal position, about 15 deg CW. Perry will test. Remounted stovepipe, shut down tel.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 4/11/16
Submitted - Monday, April 11, 2016 at 15:43:34 (MST)
Filters inspected.
Mike
Telescope Involved: 48"
Date changed: 04/07/2016
Submitted - Friday, April 08, 2016 at 05:27:03 (MST)
CO2'd primary mirror.
Mike
Telescope Involved: 48"
Date changed: 04/07/2016
Submitted - Friday, April 08, 2016 at 05:26:43 (MST)
Ton Gerl replaced the 48" dome slit batteries, adding two batteries.
wap <waynep>
Telescope Involved: 48"
Date changed: 03/08/2016
Submitted - Tuesday, March 08, 2016 at 15:14:31 (MST)
WAP+TG replaced CCD shutter power supply to address shutter failures. No failures that night, although temps were above 4C, so we'll have to see what happens when they fall well below freezing again.
Emilio Falco <falco@cfa.harvard.edu>
Telescope Involved: 1.2m
Date changed: 02/04/2016
Submitted - Friday, February 05, 2016 at 10:41:07 (MST)
FTS mirror chiller replaced with new equivalent unit, Steve, Wayne, Tom, Larry. Some teething problems, but it is now working and TG got it to communicate. Sadly, it will not operate until the pump for the building chiller is replaced, week of 02/01/16.
Emilio Falco <falco@cfa.harvard.edu>
Telescope Involved: 1.2m
Date changed: 01/27/2016
Submitted - Friday, January 29, 2016 at 16:45:54 (MST)
Steve+Larry worked with WAP. They replaced a broken axle in one of the dome bearings. Noise gone.
Emilio Falco <falco@cfa.harvard.edu>
Telescope Involved: 1.5m
Date changed: 01/25/2016
Submitted - Friday, January 29, 2016 at 16:25:00 (MST)
WAP swapped the ion pump on FAST. WAP also removed the mounting tube for the ion pump and cleaned out the crud deposited by the failed pump. After reassembly, WAP performed a full-bake of the dewar to decontaminate it.
wap <waynep>
Telescope Involved: 60"
Date changed: 11/4/2015
Submitted - Tuesday, December 01, 2015 at 15:27:43 (MST)
PB and GE tweaked the collimation. A big improvement in image quality! See the collimation notes.
Emilio Falco <falco>
Telescope Involved: 1.5m
Date changed: 11/08/2015
Submitted - Monday, November 09, 2015 at 12:56:44 (MST)
TG tweaked S hard point, under the watchful eye of EF. Final values of load cell readings: NW NE S 20.4 24.0 19.3
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 10/01/15
Submitted - Thursday, October 01, 2015 at 15:58:52 (MST)
Obtained pointing grid, to hand over to Tom Trebisky for analysis and new pointing model. Previous one was from 2010. About 210 stars for a full map, about 2 hours. Turned flex correction back on after that.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 9/30/2015
Submitted - Wednesday, September 30, 2015 at 20:53:34 (MST)
Dis- and re-mounted M1 after adjusting position. Able to collimate properly after that, finally after terrible weather start. Munged a couple of temp sensors.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 9/18/2015
Submitted - Wednesday, September 30, 2015 at 20:51:51 (MST)
TG+EF glued the tip-tilt mirror back on its stage. Used Omegabond 101 epoxy (stored in 60 fridge) that we used for new RTDs on the 1.2m M1. Accessed by removing side plate closest to the stage.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 7/16/2015
Submitted - Monday, July 20, 2015 at 11:54:20 (MST)
CO2'ed primary mirrors of both telescopes.
Mike
Telescope Involved: 60" & 48"
Date changed: 06/30/2015
Submitted - Wednesday, July 01, 2015 at 03:55:51 (MST)
EF and Gil found a cover was not closing completely. Loosening and retightening screw with nut at elbow fixed that. We also found a loose radial cw, on the S end, underneath a cover. Left it for Ted who tightened it with Larry on 6/3. Need to adjust collimation.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 6/2/15
Submitted - Monday, June 08, 2015 at 14:11:20 (MST)
Larry installed/replaced all of the casters on the 48" telescope dome upper slit. He also took the opportunity to lube the worm gear assembly which is used to open the upper slit. This should complete the 48" dome refurbishing project.
Wayne Peters <waynep>
Telescope Involved: 48"
Date changed: 4/16/15
Submitted - Thursday, April 16, 2015 at 13:38:56 (MST)
WAP completed installation, and configuration of new mirror temperature monitoring system, which employs 24 custom 100 Ohm platinum 4-wire RTDs read in a differential mode differential. System shows that old temperature monitoring system is 3 deg C high in error.
Wayne Peters <waynep>
Telescope Involved: 48"
Date changed: 4/13/15
Submitted - Thursday, April 16, 2015 at 08:29:18 (MST)
WAP baked Kep 3/10-3/12. Flats show much reduced measles. Collimation was good on 3/13, but checked with poor seeing. Adjusted radial + axial hardpoints and cws. Axial load cell readings are around 20lb and return there after slews. Great!
Falco <falco@cfa.harvard.edu>
Telescope Involved: 1.2m
Date changed: 3.1415
Submitted - Saturday, March 14, 2015 at 17:47:47 (MST)
WAP replaced single dead flat lamp (SE) with same halo type. Lasted 3 months?
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 2/27/15
Submitted - Tuesday, March 03, 2015 at 19:49:58 (MST)
Cleaned filters.
Mike
Telescope Involved: 48"
Date changed: 3/2/2015
Submitted - Tuesday, March 03, 2015 at 05:12:01 (MST)
CO2'd primary mirror
Mike
Telescope Involved: 48"
Date changed: 3/2/2015
Submitted - Tuesday, March 03, 2015 at 05:11:34 (MST)
CO2'2 primary mirror.
Mike
Telescope Involved: 60"
Date changed: 3/2/2015
Submitted - Tuesday, March 03, 2015 at 05:10:56 (MST)
C02 mirror cleaning ok
PB <perry>
Telescope Involved: 60/48
Date changed: 01/24/2015
Submitted - Saturday, January 24, 2015 at 11:57:08 (MST)
Replaced camera guide gain cable.
wap <waynep>
Telescope Involved: 60"
Date changed: 12-15-2014
Submitted - Monday, December 15, 2014 at 16:17:57 (MST)
west dome wheel was observed to be climbing off of the shaft, after the dome would not home. Removed and re-mounted the drive wheel with Larry S.
wap <waynep>
Telescope Involved: 60"
Date changed: 12/09/2014
Submitted - Tuesday, December 09, 2014 at 16:37:15 (MST)
That was WAP replaceD the flat lamps, and the model is: Hikari JCR-8191P - 10 Watt - MR11 Flood - Glass Face - 1,500 Life Hours - 6 Volt
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 12/02/2014
Submitted - Wednesday, December 03, 2014 at 17:19:57 (MST)
WAP replace the 4 dome flat lamps with new halogen lamps with reflectors. The exposure times are cut by 50% with these new lamps.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 12/02/2014
Submitted - Tuesday, December 02, 2014 at 16:43:21 (MST)
WAP and Larry finished adding counter nuts to all 40 supports for the dome wheels.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 12/01/2014
Submitted - Tuesday, December 02, 2014 at 16:42:12 (MST)
WAP replaced GE726 FLAT lamps with reflector base halogen lamps, over doubling the flat lamps output (reducing FLAT time considerably).
wap <waynep>
Telescope Involved: 1.2m
Date changed: 12/02/2014
Submitted - Tuesday, December 02, 2014 at 15:21:11 (MST)
Danny, Larry, Wayne pulled old dome tracks, replaced them. EF provided expert supervision, some muscle, drove the dome. Adjusted bearing wheels also (will add locknuts next week). Started 11/03, finished 11/07, looks and sounds much much much better. Downtime was 3 nights, not bad!
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 11/07/2014
Submitted - Friday, November 07, 2014 at 16:01:28 (MST)
Larry found the W dome wheel on the mezzanine. Bolts sheared, tube shredded. He had inspected it on 09/11/2014 and saw no problem. Lucas Macri observed the previous 3 nights; he reported nothing amiss. Larry replaced the wheel and tube, used hardened bolts.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 10/27/2014
Submitted - Saturday, November 01, 2014 at 12:55:11 (MST)
Lubed dome slit opener/closer worm gear and mechanisms.
wap <waynep>
Telescope Involved: 48"
Date changed: 10/20/14
Submitted - Monday, October 20, 2014 at 15:28:55 (MST)
C02 mirror cleaning
P Berlind <perry>
Telescope Involved: 60/48
Date changed: 10/03/2014
Submitted - Saturday, October 04, 2014 at 03:54:15 (MST)
WAP replaced all 4 flat lamps with GE 786 bulbs. They are 12W bulbs, 240 lumens. Tried Osram 6425 first, but they are only 10W, 200 lumens. Use the GEs always!
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 9/22/14
Submitted - Monday, September 22, 2014 at 17:09:48 (MST)
Tel back, working post-shutdown. WAP returned to fix the dome. Lowered the speed a little. That allows it to stop. Harvey bit to watch: I3. Dome skirt will go on next week. Collimated, barely needed a tweak.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 09/05/2014
Submitted - Saturday, September 06, 2014 at 09:26:05 (MST)
Cleaned ADC with dist water following Gabor recipe, via Mike.
E Falco <falco>
Telescope Involved: 1,5m
Date changed: 8/29/14
Submitted - Monday, September 01, 2014 at 15:01:15 (MST)
24 new temp sensors installed on M1. CO2 + dist water wash of M1. Looks much better.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 8/29/14
Submitted - Monday, September 01, 2014 at 14:58:03 (MST)
Taped up the black cowling around the TRES guide camera, at the base of the front-end, in an effort to block ambient light sources. Hopefully this will help to alleviate the position dependent fringing issue.
W Peters <waynep>
Telescope Involved: 60"
Date changed: 07/21/13
Submitted - Tuesday, July 22, 2014 at 15:17:43 (MST)
Hexapod failed again. TG found the CPU and case fans in its control PC were dead. He replaced them. The PC and hexapod are working again.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 07/07/2014
Submitted - Thursday, July 10, 2014 at 16:01:55 (MST)
Primary mirror CO2'd
M Calkins <Mike>
Telescope Involved: 60"
Date changed: 07/06/2014
Submitted - Sunday, July 06, 2014 at 23:54:56 (MST)
Filters cleaned.
M Calkins <Mike>
Telescope Involved: 48"
Date changed: 07/06/2014
Submitted - Sunday, July 06, 2014 at 23:54:25 (MST)
Primary mirror CO2'd
M Calkins <Mike>
Telescope Involved: 48"
Date changed: 07/06/2014
Submitted - Sunday, July 06, 2014 at 23:53:54 (MST)
Hexapod system failed for the last hour of night of 7/1. TG rebooted the PC twice, second time was the charm. Back in good working condition. TG reported the hexapod seems noisier than before, when it was super quiet.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 7/2/14
Submitted - Wednesday, July 02, 2014 at 15:58:25 (MST)
Neglected to mention that Danny was instrumental in the work on the CWs, providing manpower and tools!
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 7/1/14
Submitted - Wednesday, July 02, 2014 at 15:55:57 (MST)
Ted+Larry changed out all the CWs am. We measured the force w/Futek on 2 that were accessible, one 15.7, one 9.6 as expected for thick and thin CWs. Adjusted all 27 CWs to the proper positions on their shafts, so forces are nominal. Ted adjusted all axial hardpoints to 21lb, per nominal. After topbox+kepcam were on, they read 15, 22, 21lb. Recorded numbers & locations of all CWs. Came back to life in good shape!
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 7/1/14
Submitted - Tuesday, July 01, 2014 at 17:14:04 (MST)
Bypassed a relay, allowing the 60" AC unit to cool the chamber until the repairman arrives
wap <waynep>
Telescope Involved: 60"
Date changed: 6/23/2014
Submitted - Monday, June 23, 2014 at 12:06:31 (MST)
Repaired the chamber AC chiller - bad connection
wap <waynep>
Telescope Involved: 48"
Date changed: 6/16/2014
Submitted - Monday, June 23, 2014 at 12:04:02 (MST)
CO2 wash both mirrors
PB <perry>
Telescope Involved: 60/48
Date changed: 06/12/2014
Submitted - Thursday, June 12, 2014 at 14:45:18 (MST)
I fix
wap <waynep>
Telescope Involved: 48"
Date changed: 6-12-14
Submitted - Thursday, June 12, 2014 at 12:37:46 (MST)
Larry checked the small gen on the loading dock. Bad gas prevented a start. He drained it, started it. Back on monthly test schedule now.
E Falco <falco>
Telescope Involved: 1.2m+1.5m
Date changed: 6/4/14
Submitted - Wednesday, June 04, 2014 at 15:21:30 (MST)
Perry restarted TCS on mount PC at about 17:10, after domes failed. TCS ran for >1mo w/o a problem, suddenly froze as domes started.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 5/27/14
Submitted - Thursday, May 29, 2014 at 13:02:07 (MST)
Filters cleaned and inspected.
Mike
Telescope Involved: 48"
Date changed: 05/11/2014
Submitted - Monday, May 12, 2014 at 14:39:28 (MST)
Filters inspected.
Mike
Telescope Involved: 48"
Date changed: 04/12/2014
Submitted - Sunday, April 13, 2014 at 14:52:27 (MST)
WAP found blown fuse on DEC drive. Drive was making unusual noise. WAP lubed it, working well now. Also, WAP removed 1 of 5 counterweights from each post. Probably caused too much strain.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 04/07/2014
Submitted - Monday, April 07, 2014 at 17:03:45 (MST)
TG swapped MB in the mount PC, again. Had some of same troubles with telemetry. TG found settings that got it going:1 is 2F8/IRQ3, 2 is auto. After it was working, the mount PC crashed. It booted without any boards, after some reaseating of these, it came back on. Then crashed again. TG then swapped in a new power supply. That has worked for a few hours today, there is hope henceforth. But we need hot-swap spare PCs, for both tels. A challenge because the Tech 80 mod 5638-2 is no longer made, and nothing is on ebay... TG suggested a new wire to fix the guider. The NS stage was chattering, making strange noise. TG+EF swapped in a spare wire, which quieted the motor, and it's now working again. Consider more, better spare wires.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 04/06/2014
Submitted - Monday, April 07, 2014 at 17:01:38 (MST)
TG swapped MB on mount PC, which was slow and had clear problems. But that caused paralysis in the system. WAP swapped in the RA motor encoder, which had gone bad. TCS then seemed to spew out garbage rather than useful telemetry. After 2 days of wrestling with it, WAP showed wires and ports were OK. Close to COB today, back to old MB, and after WAP selected ports on TCS in the right order, everything is back to normal.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 4/4/14
Submitted - Friday, April 04, 2014 at 18:39:23 (MST)
Mirror CO2'd
Mike
Telescope Involved: 48"
Date changed: 03/06/2014
Submitted - Thursday, March 06, 2014 at 06:22:47 (MST)
Filters inspected and cleaned.
Mike
Telescope Involved: 48"
Date changed: 03/05/2014
Submitted - Thursday, March 06, 2014 at 05:17:11 (MST)
Tom Schultz of Cactus Heating & Cooling worked on mirror chiller, which had not been cooling as much as needed. He replaced a broken fan and reloaded the unit with R22. Now all we need is cold again to test.
E Falco <falco@cfa.harvard.edu>
Telescope Involved: 1.2m
Date changed: 2/21/14
Submitted - Friday, February 28, 2014 at 14:23:55 (MST)
WAP worked on the rack. I held the flashlight. He soldered a loose wire and temporarily replaced a broken resistor. After many nights of relay trips, Warren had none the night of 2/27. WAP is making the resistor permanent on 2/28/14. We may be out of the woods on this one, for now. The rack is a tangled mess of wires, poorly made boards and old components, any of which may fail anytime. The documentation is almost worse than nothing. Think new rack!
E Falco <falco@cfa.harvard.edu>
Telescope Involved: 1.5m
Date changed: 2/27/14
Submitted - Friday, February 28, 2014 at 14:17:01 (MST)
WAP worked on the rack. I held the flashlight. He soldered a loose wire and temporarily replaced a broken resistor. After many nights of relay trips, Warren had none the night of 2/27. WAP is making the resistor permanent on 2/28/14. We may be out of the woods on this one, for now. The rack is a tangled mess of wires, poorly made boards and old components, any of which may fail anytime. The documentation that is almost worse than nothing. Think new rack!
E Falco <falco@cfa.harvard.edu>
Telescope Involved: 1.5m
Date changed: 2/27/14
Submitted - Friday, February 28, 2014 at 14:16:06 (MST)
Nflwo60 computer was replaced with flwo60, a new 8-core CPU with 8GB of memory running the most recent Slackware.
Ted Groner <ted>
Telescope Involved: 1.5M
Date changed: 2/5/2014
Submitted - Tuesday, February 25, 2014 at 11:52:58 (MST)
CO2 snow blow both telescopes
PB <perry>
Telescope Involved: 60 and 48
Date changed: 02/08/2014
Submitted - Saturday, February 08, 2014 at 13:47:22 (MST)
Dec servo motor on the 60" replaced with spare.
WAP <waynep>
Telescope Involved: 60"
Date changed: 1/23/14
Submitted - Friday, January 24, 2014 at 13:44:35 (MST)
48" filters inspected.
Mike
Telescope Involved: 48"
Date changed: 12/13/2013
Submitted - Friday, December 13, 2013 at 07:01:23 (MST)
48" primary mirror CO2'd
Mike
Telescope Involved: 48"
Date changed: 12/13/2013
Submitted - Friday, December 13, 2013 at 07:00:38 (MST)
CO2'ed primary mirror.
M Calkins
Telescope Involved: 60"
Date changed: 12/11/2013
Submitted - Thursday, December 12, 2013 at 04:35:46 (MST)
Installed c-mount Custom Sci. low-pass filter (400-700nm) on guide cam lens, to filter red flashes from brakes.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 11/18/13
Submitted - Monday, November 18, 2013 at 14:31:52 (MST)
Finally collimated the tel. Used aberfit but reversing the signs of neutral tilts. That worked, not clear why the signs are wrong, but at least we know the cure.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 11/07/2013
Submitted - Friday, November 08, 2013 at 09:58:43 (MST)
48" primary mirror CO2'd
Mike
Telescope Involved: 48"
Date changed: 11/08/2013
Submitted - Friday, November 08, 2013 at 06:29:04 (MST)
60" primary mirror CO2'd
Mike
Telescope Involved: 60"
Date changed: 11/06/2013
Submitted - Wednesday, November 06, 2013 at 16:58:12 (MST)
48" filters inspected, air dusted if necessary.
Mike
Telescope Involved: 48"
Date changed: 11/05/2013
Submitted - Tuesday, November 05, 2013 at 15:58:55 (MST)
CO2'd both the 60" and 48". We were able to get all of the 60" mirror by utilizing a ladder and some ingenuity and bodily contortion.
Mike & Perry
Telescope Involved: 60" & 48"
Date changed: 10-20-2013
Submitted - Sunday, October 20, 2013 at 16:31:48 (MST)
WAP adjusted the mirror cover openers (limit switches) so relay chatter is eliminated.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 10/16/2013
Submitted - Wednesday, October 16, 2013 at 14:23:43 (MST)
Mirror cover openers are once again operational. WAP installed new DFM hardware. Some tweaks of actuators are still necessary to prevent relay chatter. WAP working on that.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 10/15/13
Submitted - Wednesday, October 16, 2013 at 13:43:08 (MST)
Filters inspected and cleaned.... The "New Z" filter had some "dirt" accumulated in one corner. All other filters looked fine.
M Calkins
Telescope Involved: 48"
Date changed: 9/13/2013
Submitted - Friday, September 13, 2013 at 06:43:44 (MST)
WAP traced ADC problem to connector to Digi. Soldered in the right wire, so it's now working.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 9/3/13
Submitted - Tuesday, September 03, 2013 at 17:33:19 (MST)
Replaced WiFi dome thermostat that was lightning damaged with new CT30 thermostat for computer wifi control.
Ted Groner <ted>
Telescope Involved: 1.2 M
Date changed: 08/28/2013
Submitted - Wednesday, August 28, 2013 at 10:09:10 (MST)
Peeled opticoat, beautiful mirror, lifted cell with Larry at the hoist. Mounted/dismounted topbox to re-adjust axials. Ended with axial loads 24, 22, 21, called it good. Radial dial indicators at 0 after setting radial hard points to 14, 15, 9 (aimed for lower but settled on these).
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 8/21/13
Submitted - Wednesday, August 21, 2013 at 15:41:23 (MST)
Bellevilles arrived. Spreader bars mounted. Mirror in cell. Adjusted axial cw's+hard points, several rounds, until mirror was at nominal 1.25in from top of cell. Radial actuators good.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 8/13/13
Submitted - Wednesday, August 21, 2013 at 15:38:39 (MST)
Still trying to get belleville washers for the load spreader bars. They are a small and unique size, which complicated the process. Waiting for a quote from a supplier...M1 awaits these for final installation.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 8/12/13
Submitted - Monday, August 12, 2013 at 15:21:59 (MST)
WAP had the base of the stovepipe cut to make space for easier intallation/removal, without affecting the ADC motors. Success!
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 8/10/13
Submitted - Monday, August 12, 2013 at 15:20:10 (MST)
8 temp sensors are epoxied to M1, top and bottom of glass. Trying to determine whether we can have spreader bars ground at the ends, to match thinner bars for old M1.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 8/7/13
Submitted - Wednesday, August 07, 2013 at 16:41:31 (MST)
Support, EF removed new M1 from box, lifted it and left it hanging about 1.5m from the ground. WAP+EF installed 4 of 8 temp sensors. Epoxy remained runny longer than advertised, takeaway was let it tack. Then it's easy, but it was late in the day because temp sensors were found late also.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 8/6/13
Submitted - Tuesday, August 06, 2013 at 16:43:02 (MST)
TG, WAP, EF, MC, support removed old M1. Brought new M1 into chamber, left overnight in its box.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 8/5/13
Submitted - Tuesday, August 06, 2013 at 16:38:22 (MST)
Broken dome W wheel (same as had been repaired a week earlier).TG and EF went up to repair Sat 6/22. Gil assisted. It had sheared its bolts and fallen out the night of 6/21 so lost that night. Had difficulty finding right bolts (size, grade) but TG succeeded, barely. Inner tube was found to be leaking after reassembly. Call to WAP led to spare inner tube. TG and Gil installed it. Working now.
Emilio <falco>
Telescope Involved: 1.5m
Date changed: 6/23/13
Submitted - Sunday, June 23, 2013 at 10:27:00 (MST)
T cleaned camera window by blowing air on 5/28. W cleaned more with CO2 on 5/29. E finished by daubing lightly with lens cleaner and tool to reach in, with final CO2 by W. Window looks pristine now.
E Falco <falco>
Telescope Involved: PAIRITEL
Date changed: 5/29/2013
Submitted - Wednesday, May 29, 2013 at 16:57:49 (MST)
W+E tightened ADC loose screw again, applied loctite this time. Reinstalling the baffle caused a connector to come unplugged, disabling ADC until 5/28.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 5/24/2013
Submitted - Wednesday, May 29, 2013 at 16:55:23 (MST)
east mirror cover disabled.
WP <waynep>
Telescope Involved: 48"
Date changed: 5/14/13
Submitted - Monday, May 20, 2013 at 16:13:37 (MST)
ADC set screw tightened
w p <waynep>
Telescope Involved: 60"
Date changed: 5/13/13
Submitted - Monday, May 20, 2013 at 16:12:40 (MST)
KeplerCam fill times adjusted to 5:30pm and 7:15am
wap <waynep>
Telescope Involved: 48"
Date changed: 4/18/13
Submitted - Thursday, April 18, 2013 at 16:40:19 (MST)
Fuse replaced in the dome fused disconnect. I will need to replace the controller tomorrow. There is no speed adjustment. The dome rotates at full speed. There may be some "washing machine action" experienced tonight..
wap <waynep>
Telescope Involved: 60"
Date changed: 4/18/13
Submitted - Thursday, April 18, 2013 at 16:37:27 (MST)
T Groner and D Gallup removed the 1.5M shroud, exposing the ADC. The A motor gear set screw was found to have loosened. Screw was tightened, and shroud replaced.
Ted Groner <ted>
Telescope Involved: 1.5 M
Date changed: 02/07/2013
Submitted - Thursday, February 07, 2013 at 11:35:43 (MST)
WP+TG+Don worked on dome. Repaired mechanical failure of one piston and replaced voltage reg so it all works again.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 01/25/13
Submitted - Saturday, January 26, 2013 at 17:07:14 (MST)
TG updated Keplercam CCD bias levels. Amps 3 and 4 were showing levels about 20 counts higher than 1 and 2. Now they are back to all about the same.
Emilio <falco>
Telescope Involved: 1.2m
Date changed: 01/11/13
Submitted - Sunday, January 13, 2013 at 13:58:26 (MST)
Gabor Furesz replaced the two stages for the Tres projector arm. G. Furesz and T. Groner reworked the software and aliases to produce much faster homing and comparison routines.
Ted Groner <ted>
Telescope Involved: 1.5M
Date changed: 01/10/2013
Submitted - Thursday, January 10, 2013 at 11:58:32 (MST)
W Peters changed the wireless dome relay control card. Function tested by T Groner after replacement.
T Groner <ted>
Telescope Involved: 1.2M
Date changed: 12/26/2012
Submitted - Wednesday, December 26, 2012 at 13:20:26 (MST)
EF and TG cleaned the ADC with alcohol and gentle daubs with kimwipes. It looked filthy, looks much better now. Warren reported a decrease in throughput with FAST, which should be better now.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 11/29/2012
Submitted - Thursday, November 29, 2012 at 16:01:05 (MST)
Lost night of 9/18 to stuck mirror covers. WAP fixed on 9/19 in spite of his cold. A connector was not seated properly...
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 9/19/12
Submitted - Wednesday, September 19, 2012 at 16:31:13 (MST)
ugriz filters returned from Custom Scientific, looking nice and clean
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 08/15/12
Submitted - Wednesday, August 15, 2012 at 14:50:54 (MST)
ugriz filters were very dirty after the rain event. They were sent to Custom Scientific for evaluation and recovery if possible. Evaluation is still underway as of 08/10.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 07/19/12
Submitted - Friday, August 10, 2012 at 17:01:20 (MST)
WAP and TG chased the failure due to the rain event in one of the topbox motor controller cards and managed to revive the card using parts from a spare card. As of today, the topbox is back in working condition.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 07/25/2012
Submitted - Wednesday, July 25, 2012 at 16:05:38 (MST)
Dome slit top remained open after observations. PB left the 1.5m at 4 am and didn't notice in the dark. It rained during the day and it wasn't noticed until about 3pm when Ale and Warren came down to check on the generator, which was not generating after power was lost. Wayne came up and closed the slit with help from a battery charger. Due to clouds, the batteries had not recharged and caused the failure.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 07/15/2012
Submitted - Wednesday, July 25, 2012 at 16:02:54 (MST)
WAP repaired leaking coolant line on the cell. Back to cooled mirror.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 05/29/12
Submitted - Wednesday, May 30, 2012 at 11:59:33 (MST)
Filters inspected.
M Calkins <Mike>
Telescope Involved: 48"
Date changed: 05/03/2012
Submitted - Thursday, May 03, 2012 at 16:43:25 (MST)
Filters inspected and cleaned.
M Calkins
Telescope Involved: 48"
Date changed: 01/27/2012
Submitted - Friday, January 27, 2012 at 19:13:00 (MST)
CO2ed the mirror, no noticeable improvement.
M Calkins
Telescope Involved: PAIRITEL
Date changed: 01/27/2012
Submitted - Friday, January 27, 2012 at 18:51:24 (MST)
CO2ed the mirror, no noticeable change.
M Calkins
Telescope Involved: 48"
Date changed: 01/27/2012
Submitted - Friday, January 27, 2012 at 18:50:48 (MST)
CO2ed the mirror... Noticeable improvement. Perhaps recent dome maintenance were the culprit of dust build-up?
M Calkins
Telescope Involved: 60"
Date changed: 01/27/2012
Submitted - Friday, January 27, 2012 at 18:49:58 (MST)
Replaced guider focus motor with the same Vextra model as the original. The replacement had very minor differences compared to the original motor. It required turning 2 holes in the chassis into through holes and cutting off 1/2 of its axle. Hardest part was removal of old motor. Set screws in its coupling to the shaft were impossible to remove. Loctite remover did nothing. Had to drill out one of the screws. Reassembly went well. Started 12/07, on the sky the night of 12/08. Guider worked well that night. Assistance provided by Marc Lacasse and Rodger Harris. Wayne, Ted, Emilio.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 12/08/11
Submitted - Friday, December 09, 2011 at 11:13:36 (MST)
West 1.5m dome motor failed. Motor was replaced by P Berlind, T Welsh and T Groner.
T Groner <tedg>
Telescope Involved: 1.5m
Date changed: 11/09/2011
Submitted - Wednesday, November 09, 2011 at 09:31:51 (MST)
WP baked Keplercam.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 08/29/2011
Submitted - Monday, September 26, 2011 at 11:21:54 (MST)
Cleaned filters, CO2'd mirror, cleaned flat lamps (which were horribly filthy!). Flat counts reported to have increased by over 3,000.
Mike
Telescope Involved: 48"
Date changed: 07/28/2011
Submitted - Wednesday, August 03, 2011 at 22:30:39 (MST)
Filters inspected
Mike
Telescope Involved: 48"
Date changed: 6/27/2011
Submitted - Wednesday, June 29, 2011 at 05:28:50 (MST)
Filters inspected.
Mike
Telescope Involved: 48"
Date changed: 04/27/2011
Submitted - Thursday, April 28, 2011 at 05:14:50 (MST)
Filters inspected
Mike
Telescope Involved: 48"
Date changed: 03/11/2011
Submitted - Saturday, March 12, 2011 at 05:44:03 (MST)
Filters inspected.
Mike
Telescope Involved: 48"
Date changed: 02/18/2011
Submitted - Saturday, February 19, 2011 at 06:01:24 (MST)
Enabled the ADC system now for both FAST and TRES use. The system comes on when the Realtime system is started.
Ted Groner <ted>
Telescope Involved: 1.5 M
Date changed: 1/28/2011
Submitted - Saturday, January 29, 2011 at 18:17:46 (MST)
On 12/15/10, TG and I obtained a new pointing grid. Dave Harvey processed the data and sent us a new flexure corrections file on 01/03/11. Proper tests on 01/07/11 showed pointing is greatly improved, errors < 11 asec in DEC and <14 asec in RA all over the sky, except 17 and 22 asec in RA at -3h and -4h, respectively.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 01/05/11
Submitted - Friday, January 21, 2011 at 13:44:05 (MST)
Filters cleaned/inspected.
Mike
Telescope Involved: 1.2
Date changed: 11/02/2010
Submitted - Friday, November 12, 2010 at 06:20:49 (MST)
TG added commands, WAP connected relays in the rack to: (1) open/close the mirror covers (40 sec) and (2) turn on/off the flat lamps. Telshell commands tele coversup (dn) and tele incandon (off).
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 9/21/2010
Submitted - Wednesday, September 22, 2010 at 11:41:51 (MST)
WAP installed board, TG updated software to open/close dome slit in 165 seconds. Exercised it several times, success! Commands are sent from Telshell.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 09/03/2010
Submitted - Friday, September 03, 2010 at 17:06:56 (MST)
WAP and Steve Brown completed the new, much improved system to open the dome slit. Manual mode for now, computer mode soon.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 09/01/2010
Submitted - Thursday, September 02, 2010 at 12:38:20 (MST)
No collimation necessary! Dome home tab had been removed by Tom Welsh, thinking it unnecessary. Re-mounted 09/02.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 09/01/2010
Submitted - Thursday, September 02, 2010 at 12:35:05 (MST)
PB, TG and EF collimated as weather had cleared. It took little time, got a good collimation.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 09/01/2010
Submitted - Thursday, September 02, 2010 at 12:33:11 (MST)
Joe Zajac sent newly-coated (hard, not sol-gel) optics, 2nd set. Gabor Furesz sent the can with the old grimy optics, arrived 08/31. Replaced old w/newly-coated. Very tight fit, had to use heat tape to expand can a little, which allowed the optics to fit.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 09/01/2010
Submitted - Thursday, September 02, 2010 at 12:31:32 (MST)
Re-aluminized M2 early Aug at Sunnyside. Gil applied CO2 and then washed M1 (daubed w/distilled water) mid-Aug, looks much better.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 09/01/2010
Submitted - Thursday, September 02, 2010 at 12:28:56 (MST)
48" mirror "cleaned" with CO2.
Mike
Telescope Involved: 48"
Date changed: 06/06/2010
Submitted - Sunday, June 06, 2010 at 17:36:46 (MST)
48" filters inspected/cleaned.
Mike
Telescope Involved: 48"
Date changed: 06/06/2010
Submitted - Sunday, June 06, 2010 at 17:35:48 (MST)
48" filters inspected...
Mike <Mike>
Telescope Involved: 48"
Date changed: 05/09/2010
Submitted - Sunday, May 09, 2010 at 23:34:04 (MST)
WAP was finally able to get a working switch for the dome slit. He installed it with Steve Brown up near the top of the slit, which required the lift. PAIRITEL is back to fully automatic operation.
E Falco <falco>
Telescope Involved: PAIRITEL
Date changed: 4/27/2010
Submitted - Wednesday, April 28, 2010 at 12:38:14 (MST)
60" snow blow. Fresh tank in use. Also snow blow 3/20
PB <perry>
Telescope Involved: 60"
Date changed: 04/23/2009
Submitted - Friday, April 23, 2010 at 13:26:19 (MST)
Primary mirror CO2'd, filters inspected.
Mike
Telescope Involved: 48"
Date changed: 03/28/2010
Submitted - Sunday, March 28, 2010 at 19:03:14 (MST)
Inspected PAIRITEL mirror. Looked good. No need for CO2 blowing.
Mike
Telescope Involved: PAIRITEL
Date changed: 3/14/2010
Submitted - Tuesday, March 23, 2010 at 22:00:43 (MST)
Used CO2 to clean mirror surface.
Mike
Telescope Involved: 48"
Date changed: 2/28/2010
Submitted - Sunday, February 28, 2010 at 22:42:06 (MST)
Filters cleaned/inspected
Mike
Telescope Involved: 48"
Date changed: 02/28/2010
Submitted - Sunday, February 28, 2010 at 22:41:17 (MST)
Filters inspected
Mike
Telescope Involved: 48"
Date changed: 01/31/2010
Submitted - Monday, February 01, 2010 at 16:25:28 (MST)
60" CO2 snow blow. A few drip marks on the primary.
PB <perry>
Telescope Involved: 60"
Date changed: 01/17/10
Submitted - Monday, January 18, 2010 at 01:29:15 (MST)
CO2 snow cleaning of primary.
Mike
Telescope Involved: 48"
Date changed: 01/10/2010
Submitted - Sunday, January 10, 2010 at 15:30:34 (MST)
Filters cleaned/inspected.
Mike
Telescope Involved: 48"

Submitted - Sunday, January 10, 2010 at 15:29:34 (MST)
CO2 snow-blow cleaning of 60" primary. ok.
PB <perry>
Telescope Involved: 60"
Date changed: 12/14/09
Submitted - Monday, December 14, 2009 at 15:15:55 (MST)
MC applied CO2 snow on 11/14. EF and WP cleaned M2 and M1 with Edmunds blue, Kimwipes and cotton. Both look much better. Also wiped filthy covers, using Kimwipes and water.
E Falco <falco>
Telescope Involved: PAIRITEL
Date changed: 11/16/09
Submitted - Tuesday, November 17, 2009 at 13:12:14 (MST)
CO2'd primary mirror...
MC <mike>
Telescope Involved: 48"
Date changed: 11/15/2009
Submitted - Monday, November 16, 2009 at 02:19:35 (MST)
Cleaned/inspected filters.
MC <mike>
Telescope Involved: 48"
Date changed: 11/14/2009
Submitted - Saturday, November 14, 2009 at 17:41:47 (MST)
CO2'd the primary mirror. Removed one menacing build-up/spot, but mirror could use a more thorough cleaning.
MC <mike>
Telescope Involved: PAIRITEL
Date changed: 11/14/2009
Submitted - Saturday, November 14, 2009 at 17:41:13 (MST)
60" primary CO2 snow blow. Mirror looks very clean. Also snow blow 9/24/09
PB <perry>
Telescope Involved: 60"
Date changed: 10/28/09
Submitted - Wednesday, October 28, 2009 at 15:13:33 (MST)
With TG, EF replaced obviously blown fuse on dome slit opener, re-engaged gear. It operates normally after last night's failure. The cause of the fuse blowing remains unknown.
Emilio Falco <falco>
Telescope Involved: 1.5m
Date changed: 10/16/09
Submitted - Friday, October 16, 2009 at 14:25:28 (MST)
Filters cleaned/inspected...
Mike Calkins <Mike>
Telescope Involved: 48"
Date changed: 09/04/2009
Submitted - Friday, September 04, 2009 at 22:43:29 (MST)
Filters inspected/cleaned
Mike <mike>
Telescope Involved: 48"
Date changed: 07/29/2009
Submitted - Wednesday, July 29, 2009 at 05:03:56 (MST)
The rack stopped powering up. There was a nearby lightning hit, but nothing obviously burned on 07/09. WAP & TG got it working bypassing some components. On 07/10/09, WAP found solder problems in a board, which he corrected and which brought the rack back to life.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 07/10/09
Submitted - Monday, July 13, 2009 at 15:42:38 (MST)
Filters inspected/cleaned.
Mike Calkins <mike>
Telescope Involved: 48"
Date changed: 07/03/2009
Submitted - Friday, July 03, 2009 at 20:14:19 (MST)
Filters inspected/cleaned
Mike Calkins <Mike>
Telescope Involved: 48"
Date changed: 06/07/2009
Submitted - Saturday, June 06, 2009 at 21:47:38 (MST)
Filters cleaned/inspected
Calkins <mike>
Telescope Involved: 48"
Date changed: 04/29/2009
Submitted - Wednesday, April 29, 2009 at 18:32:47 (MST)
TG, PB and Dave Harvey worked on a new pointing model. Dave fixed problem that prevented access to catalogs. Dave changed servo offset parameter, so when tel reaches s/w limit, s/w doesn't drive it beyond limit. The upshot was improved pointing except beyond DEC +65 and below -15 deg where about 1' of DEC error remains. The remaining error may not be a flexure but movement in the primary, so the pointing model doesn't correct it. We'll test this...
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 04/30/09
Submitted - Thursday, April 09, 2009 at 15:25:57 (MST)
WAP blew air on Kepcam window, donuts are gone.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 04/09/09
Submitted - Thursday, April 09, 2009 at 15:17:27 (MST)
EF and Warren Brown cleaned the corrector. It was dusty and looked milky and with dried spots on the top surface of the glass. We disassembled the unit to clean all surfaces. Blowing helped remove some dust. An isopropyl alcohol solution for optics helped a bit more. We squirted it on the surfaces and gently drew lens cleaning tissues over them. But the result was visually very disappointing. The quality remained milky and we could not remove the dried drops. We'll take throughput measurements asap.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 3/20/09
Submitted - Thursday, March 19, 2009 at 17:35:52 (MST)
The dome started making ugly noises and it eventually failed. WAP worked with Rodger+Robert to replace the gear+worm at the top of the dome. The dome now opens/closes normally, although some tweaks are still necessary.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 03/03/09
Submitted - Thursday, March 05, 2009 at 14:17:09 (MST)
Filters cleaned/inspected.
Mike Calkins <mike>
Telescope Involved: 48"
Date changed: 03/02/2009
Submitted - Monday, March 02, 2009 at 17:15:53 (MST)
Checked echelle focus: mean shift = -0.65 at micrometer setting 0.155; T = 49F. OK. Left setting unchabged
robert stefanik <rstefanik>
Telescope Involved: 1.5m
Date changed: 11 FEB 2009
Submitted - Wednesday, February 11, 2009 at 17:09:35 (MST)
Filters inspected/cleaned.
Mike Calkins <Mike>
Telescope Involved: 48"
Date changed: 02/06/2009
Submitted - Friday, February 06, 2009 at 15:28:44 (MST)
Filters cleaned.....
Mike Cakjubs
Telescope Involved: 48"
Date changed: 01/02/2009
Submitted - Friday, January 02, 2009 at 21:31:02 (MST)
Filters cleaned.
Mike Calkins
Telescope Involved: 1.2m
Date changed: 12/22/2008
Submitted - Monday, December 22, 2008 at 18:57:29 (MST)
Power supply on hexapod PC died 12/16. Replacement came in 12/17 but hexapod would not move. TG and WAP worked on it, found+fixed broken connector, and determined which pot to turn to get required 24V to hexapod. So hexapod is back in action as of 12/18. Ordered a spare power supply for the next time.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 12/18/08
Submitted - Thursday, December 18, 2008 at 17:36:04 (MST)
FAST and TRES dewars warmed up, apparently because of missed fills during the weekend of 10/18-19. TRES cooled by itself after fill. FAST was pumped during the day on 10/20. It reached 5e-6 Torr but the vacion kept "rolling." FAST was mounted 10/20. It performed well that night, so it seems to have survived without further action.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 10/20/2008
Submitted - Tuesday, October 21, 2008 at 14:42:00 (MST)
Mark Ordway and Kevin Bennett came from Cambridge, worked 10/20-21 on focus mechanism. Diagnostic: low torque on motor, no apparent problem with bearings. As soon as possible, we will replace the motor and its controller, which are decades old. Will consult with Bob Fata and Tom Gauron. Ted, Robert & Emilio assisted.
E Falco <falco>
Telescope Involved: 60
Date changed: 10/21/2008
Submitted - Tuesday, October 21, 2008 at 14:32:00 (MST)
Replaced failed UPS with a Ferrups 18kVA, replaced batteries. Stark electric removed the old, installed the new.
E Falco <falco>
Telescope Involved: all
Date changed: 10/21/2008
Submitted - Tuesday, October 21, 2008 at 14:28:20 (MST)
Filters cleaned....
Calkins <mike>
Telescope Involved: 1.2m
Date changed: 10/04/2008
Submitted - Saturday, October 04, 2008 at 18:22:10 (MST)
Issue was poor image quality likely due to M1 support misadjustment. WP dropped kepcam+topbox to adjust axial cws. EF and WP adjusted these AND axial hardpoints until load cells all read around 20lb for axial readings. Switched to radial readings, found 160lb readings, adjusted radial cws 4-6 to hang loose; also tweaked hardpoints until readings were around 20lb. Went on the sky, found collimation was good, ran kfindfwhm, kdodefoc and aberfit and ended up with a good collimation that held at large HA and ZA by 7:30pm. Then it rained all night...
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 09/25/08
Submitted - Friday, September 26, 2008 at 14:55:29 (MST)
Cleaned the corrector with help from Perry & Ted. It had dust on it and several spots as if some water had fallen on it. Daubed the surface per the instructions with Kimwipe tissues and isopropyl alcohol. Appearance improved significantly.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 09/10/08
Submitted - Thursday, September 11, 2008 at 15:27:49 (MST)
Swapped the Declination motor/encoder do to telescope runaways.
WAP <waynep>
Telescope Involved: 1.2m
Date changed: 9/5/08
Submitted - Monday, September 08, 2008 at 17:07:57 (MST)
Read reflectivity below. The facility was Sunnyside, once again.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: Aug 08
Submitted - Tuesday, September 02, 2008 at 15:28:46 (MST)
We re-aluminized the mirrors on 08/15/08. The process was successful, with good adherence and reflectiviy on the test strips. The mirror surface continues to degrade.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: Aug 2008
Submitted - Tuesday, September 02, 2008 at 15:27:00 (MST)
David Marcus, of Custom Scientific, refurbished the U filter, which had developed a leak of CuSO4. RH tested for leaks and after about 1 week, mounted the filter in the wheel. It has not leaked as of 02/04/08.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: Jan 2008
Submitted - Monday, February 04, 2008 at 16:15:36 (MST)
WAP baked FAST after an inadvertent warm-up. It looks good again, throughput has recovered.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: mid-Jan 2008
Submitted - Monday, February 04, 2008 at 16:12:48 (MST)
I replaced the chamber FOM used for remote guide camera gain. Remote camera gain now works.
Wayne P. <waynep>
Telescope Involved: 48"
Date changed: 1/15/2008
Submitted - Tuesday, January 15, 2008 at 11:28:24 (MST)
We received the replacement U filter on 06/06. Today, RH and EF filled the housing with CuSO4. What looked like a plug in the fill hole to be drilled out was a plastic screw with a hex head. As RH tried to drill, it fell inside the housing, where it will henceforth reside. It sticks to a corner and is about 2x3 mm, so we hope it will not be a problem.
Emilio Falco <falco>
Telescope Involved: 1.2m
Date changed: 06/08/07
Submitted - Friday, June 08, 2007 at 14:56:49 (MST)
FAST warmed partially on 04/18/07. The result was a "crackling" pattern visible in flats with ridges up to 10% above the average. The pattern was visible to the eye on the surface of the CCD. Consultation with D. Ouellette, M. Lesser and J. Geary yielded a consensus of deposits on the CCD that could be removed by baking or eventually cleaning. FAST once again warmed sometime over the weekend of 05/05/07. It was found at about room temp the morning of 05/07. Wayne replaced the ion pump, which had plenty of crud inside, and cleaned significant crud from the intake pipe. He conducted a sucessful bake of FAST that night. He nurtured it overnight and into the morning of 05/08. FAST was producing useable data by about noon, when Wayne finally got to go home after about 25 hours of work. The chip was nice and cold by the start of observations the night of 05/08. The first thing that was clear was that the "crackling" pattern was gone. Looking at a ratio of flats from 05/08 and 02/11, the blue sensitivity seems to have improved quite a bit (peaking at 1.3) and decreased some in the red (down to 0.9). It will be good to see what Warren's throughput measurements yield.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 05/08
Submitted - Wednesday, May 09, 2007 at 12:26:34 (MST)
Fast3 ccd replaced fast2 as of June 20th. Reworked most of fast system for this change.
Ted Groner <ted>
Telescope Involved: 1.5M
Date changed: 06/20
Submitted - Monday, July 03, 2006 at 15:35:50 (MST)
WAP replaced the connector for the gain control of the guide camera. It failed the night before.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 03/08/2006
Submitted - Wednesday, March 08, 2006 at 16:41:21 (MST)
Worked on dome drive: replaced west tire, moved gearbox so that tire makes full contact with drive surface. Added air to east tire.
R. Hutchins
Telescope Involved: 60"
Date changed: 2/17/06
Submitted - Friday, February 17, 2006 at 16:48:29 (MST)
Swapped the batteries and "SHUT" switch in the dome slit control box.
Wayne <waynep>
Telescope Involved: 60"
Date changed: 2-7-06
Submitted - Tuesday, February 07, 2006 at 15:00:13 (MST)
Completed cell chiller repairs. Hoses were replaced and flushed with clear water. Fouled heat exchangers were replaced. Coolant relaced with aluminum friendly water/glycol mix. System has a high flow rate @38 PSI.
R. Hutchins
Telescope Involved: 48"
Date changed: 2/1/06
Submitted - Monday, February 06, 2006 at 11:12:35 (MST)
Replaced RA drive motor. Old motor was failing and caused RA drive to lock up.
R. Hutchins
Telescope Involved: 60"
Date changed: 1/18/06
Submitted - Monday, February 06, 2006 at 11:06:58 (MST)
Aligned and repaired flexible RA drive coupling and alligned worm/worm gear. RA motor repairs had overstressed the flexible coupling and caused it to mis-align.
R. Hutchins
Telescope Involved: 60"
Date changed: 1/26/06
Submitted - Monday, February 06, 2006 at 11:04:09 (MST)
RH and support worked on the dome, which had been cranky. An insufficient number of the 32 axial rollers were making contact with the base of the dome. After adding shims, about 70% of the rollers make contact most of the time, which is near the design goal.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 09/21-10/05/05
Submitted - Friday, October 21, 2005 at 14:25:26 (MST)
WAP discovered that the rash of dome problems have been mechanical, not electrical. The encoder measuring wheel would not contact the dome in certain points of rotation. The radial rollers which are designed to constrain the lateral motion of the dome were at the extent of their adjustment slots. WAP loosened, adjusted and tightened some, informed EF and BH and spoke with Observadome about the correct adjustment procedure. WAP and BH adjusted the rest of the rollers and the dome is positioning correctly.
WAP <waynep>
Telescope Involved: 60"
Date changed: 9-13 thr 9-15
Submitted - Thursday, September 15, 2005 at 15:07:39 (MST)
WAP drilled & tapped blowhole into Keplercam. It is now nicely aerated through a hollowed-out screw that replaces the old piece of brass tubing.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 06 Sep 05
Submitted - Tuesday, September 06, 2005 at 15:49:18 (MST)
We shipped FAST to John Geary and Steve Amato during shutdown. They checked it, "repaired a bad driver on one of the parallel phases," and returned it in time to start observations 01 Sep 2005.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: Aug 2005
Submitted - Thursday, September 01, 2005 at 13:37:33 (MST)
WAP and PB worked with EF on the collimation below. As usual, all were involved including PB and MC in general dismantling and reassembly of the telescope after re-aluminization.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: Aug 2005
Submitted - Thursday, September 01, 2005 at 13:35:11 (MST)
Collimated the 1.5m. This time, was more difficult than usual, as the primary was not staying put. A lot of fiddling with radial and axial counter-weights finally yielded good results. Measured motions with dial indicators. WAP adjusted the axial counter-weights on the N side to reduce lever arm. With the primary in the cell, had to RTV the S spring restraint puck, which got pulled, and had to RTV the W hard point puck, which got pulled. All 3 of the latter had been RTVed after aluminization, but W didn't hold. Bot of primary ended up about 1.3 in above bot of cell (measurements vary between 1.25 and 1.4).
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 31 Aug 05
Submitted - Thursday, September 01, 2005 at 13:32:12 (MST)
Re-aluminized the primary and secondary at Sunnyside. The Al passed the tape test. Reflectivity measurements are good and are posted.
E Falco <falco>
Telescope Involved: 1.5m
Date changed: 19 Aug 05
Submitted - Thursday, September 01, 2005 at 13:25:33 (MST)
TG & EF tightened fittings to SW heat exchanger to reduce glycol leak. Also fitted a diaper around leaky fitting. Anxiously awaiting RH's return from Turkey...
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 05/05/05
Submitted - Thursday, May 05, 2005 at 14:27:40 (MST)
TG & RH lowered minicam to find dust bunnies in the corrector that caused donuts on images. There were obvious specks of dust on the top of the corrector, and between the top and the next element. Probably jostled when we put in the 4shooter mask. EF & TG blew off dust from the top and removed it then blew off other dust on the 2nd element. Images taken after that show no donuts, so success.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 04/28/05
Submitted - Thursday, April 28, 2005 at 14:32:08 (MST)
WAP & TG warmed minicam up to 30C, back down to outgas the crud on the chip. It diminished from about a 10% increase in counts to about 2%, so success. The process took all day and into the night and early morning of 04/28, so no data unfortunately.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 04/27/05
Submitted - Thursday, April 28, 2005 at 14:28:44 (MST)
test
ted <ted>


Submitted - Thursday, February 03, 2005 at 11:38:54 (MST)
Test
ted <ted>
Telescope Involved: none
Date changed: none
Submitted - Wednesday, February 02, 2005 at 13:41:14 (MST)
WAP baked minicam 01/24-25, succeeded in removing thick crud on chip 2. Some structure remains, but at acceptable levels. RH checked every actuator and adjusted those that were out of spec for balance. Two required adjustments to avoid hitting the cell. Axial load cell readings improved, from a max difference of 10lb to 5lb. Weather prevented confirmation of improvement on images. Will attempt asap.
E Falco <falco>
Telescope Involved: 1.2m
Date changed: 01/25/2005
Submitted - Monday, January 31, 2005 at 13:27:21 (MST)
Replace filter wheel controller with hot spare. Put cables three and four back in original positions. Tie-wrapped fiber-serial connector. Screwed down motor cable without backshell.
Ted <ted>
Telescope Involved: 48
Date changed: 11-03-04
Submitted - Wednesday, November 03, 2004 at 15:39:40 (MST)
WAP, TG, EF came up to look at filter wheel problems. Could not reproduce failure. WAP swapped controller & cable 3 & 4.
E Falco <falco>
Telescope Involved: 48
Date changed: 10/30/04
Submitted - Saturday, October 30, 2004 at 16:31:24 (MST)
RH repaired the dry-air compressor, which is essential for Minicam. He obtained parts from an old compressor.
Emilio Falco <falco>
Telescope Involved: 60/48
Date changed: OCt 04
Submitted - Wednesday, October 13, 2004 at 16:13:23 (MST)
RH re-glued to the side of the primary the support point for one of the spreader bars (roughly on E side). This was the same bar that came loose during re-aluminization. The top holder was re-glued when we mounted the primary, but the bottom one came loose very quickly and needed to be re-glued (with epoxy).
Emilio Falco <falco>
Telescope Involved: 48
Date changed: 20 Sep 04
Submitted - Thursday, October 07, 2004 at 11:17:42 (MST)
WP UV-flooded FAST to start the new observing season.
Emilio Falco <falco>
Telescope Involved: 60
Date changed: Sep 04
Submitted - Thursday, October 07, 2004 at 11:11:47 (MST)
Moved the slit viewer TV head out about 1 mm
Dave Latham <latham>
Telescope Involved: 60" echelle
Date changed: 27 June 2004
Submitted - Sunday, June 27, 2004 at 13:52:36 (MST)
Wayne replaced the flat lamp power supply. V flats required almost doubling exposure times. After removing 1 of 2 diffusers on all 4 lamps, we were able to obtain the same flat levels as with the old power supply.The variations of the flat levels throughout 1 night are now < 1%. With the old supply, we saw up to 15% variations!
Emilio Falco <falco>
Telescope Involved: 1.2m
Date changed: 06 May 2004
Submitted - Thursday, May 06, 2004 at 18:14:44 (MST)
echelle oil-coupling redone. echlele head removed, broken video wire fixed.
PB <perry>
Telescope Involved: 60"
Date changed: week of May 5
Submitted - Thursday, May 06, 2004 at 16:12:53 (MST)
WP and EF with RH and TG help at crucial extraction and re-insertion points of outer box, swapped cold heads in stelircam. Used refurb that ASz sent out. Stelircam appears to be on its way back as of 03/13/03, although A3 temp sensor was shorted. Its leads were cut for now, we'll need to reopen the box after the 17-20 Mar IR run.
Emilio Falco <falco>
Telescope Involved: 1.2m
Date changed: 03/12/03
Submitted - Thursday, March 13, 2003 at 16:21:05 (MST)
Scraped the permanent marker writing off of the FAST tv monitor again.
PB <perry>
Telescope Involved: 60"
Date changed: 2/1/03
Submitted - Saturday, February 01, 2003 at 18:19:33 (MST)
After confirming the decline in the FAST sensitivity, we had Mike Lesser clean the CCD and check for vacuum leaks. That removed some bad features seen in flats. The sensitivity is now about 80% of what it was 2 years ago. We are proceeding with a replacement, starting with FAST 2. John Geary is working on it in Cambridge.
E. Falco <falco>
Telescope Involved: 1.5m
Date changed: 12/02
Submitted - Tuesday, January 07, 2003 at 15:27:17 (MST)
EF+RH removed corrector. It looked very dirty to the eye. EF attempted to clean it by squirting Edmund alcohol for glass with the assembly upside down, and letting the alcohol drain out. Daubed very lightly on some remaining drops, allowed the rest to air-dry in the clean room. There was no visibly great improvement. Investigated re-coating or hard-coating and assembling the twin set of optics in hand at Cambridge, with DGF. He suggested ASz might be able to do it. ASz viewed the corrector on 11/18/02, and agreed to find plans and investigate further once back in Cambridge.
E. Falco <falco>
Telescope Involved: 1.5m
Date changed: 11/06/02
Submitted - Tuesday, November 19, 2002 at 11:46:40 (MST)
Changed to faster dual processors on flwo60. Removed fast keywords INSFOC and POSANGLE and replaced with FASTFOC and ROTANGLE, which are automatically filled. Changed this system Changelog to no longer publish email addresses due to bloody spambots!
Ted Groner <ted>
Telescope Involved: All
Date changed: 11/13/02
Submitted - Wednesday, November 13, 2002 at 13:48:12 (MST)
WAP swapped FAST stepper motor cables (focus<->sky/comp). That eliminated the problem of lost focus at restart of rtshell. The underlying cause remains unclear.
E. Falco <falco>
Telescope Involved: 1.5m
Date changed: 11/08/02
Submitted - Wednesday, November 13, 2002 at 12:26:43 (MST)
RH looked at diaphragms on the vacuum pump, which was pumping much more slowly than normal. They were clean. RH cleaned out Penning gauge for the vacuum pump. He found a <1cm piece of (probably) neoprene inside it and found that 2 O-rings were apparently not installed properly. RH corrected these faults. However, on 11/12/02 WAP started IRCAM on the pump and found it was still not pumping at a normal rate. Still to be tried: bake the pump, open a stage where there is a filtering mesh to look for other debris.
E. Falco <falco>
Telescope Involved: 1.2m, 1.5m
Date changed: 11/08/02
Submitted - Wednesday, November 13, 2002 at 12:23:18 (MST)
WAP baked out 4sh after we noticed degraded flats. Procedure: vacuum, 50C for several hours (overnight desired but insufficient time during eng run), then O2 1atm 50C UV on, overnight. Flats are much flatter now.
E. Falco <falco>
Telescope Involved: 1.2m
Date changed: 11/05/02
Submitted - Wednesday, November 13, 2002 at 10:00:22 (MST)
Reduced flow rate in AFOE fill tube by inserting a .1" aperture plate.
R. Hutchins
Telescope Involved: 60"

Submitted - Friday, April 19, 2002 at 11:37:26 (MST)
Replaced AFOE N2 line with a stainless steel hose and a larger diameter fill tube.
R. Hutchins
Telescope Involved: 60"

Submitted - Thursday, April 11, 2002 at 09:07:20 (MST)
Dome A/C condenser fan replaced.
R. Hutchins
Telescope Involved: 48"

Submitted - Tuesday, March 26, 2002 at 15:18:10 (MST)
Roger replaced east dome tire.
PB <perry>
Telescope Involved: 60"
Date changed: 1/17/01
Submitted - Saturday, January 19, 2002 at 02:46:07 (MST)
Wayne and Roger replaced west dome tire.
PB <flwo60>
Telescope Involved: 60"
Date changed: 1/2/02
Submitted - Thursday, January 03, 2002 at 00:05:13 (MST)
Replaced lower dome slit cable.
R. Hutchins
Telescope Involved: 48

Submitted - Friday, December 07, 2001 at 16:24:14 (MST)
Steve Amato
PB <perry>
Telescope Involved: 60"
Date changed: 11/14/01
Submitted - Wednesday, November 14, 2001 at 23:17:20 (MST)
today we pulled and cleaned the echelle slit plate.
PB <perry>
Telescope Involved: 60"
Date changed: 11/14/01
Submitted - Wednesday, November 14, 2001 at 23:11:15 (MST)
we are now using a new FIFO that Steve Amado sent out
PB <perry>
Telescope Involved: 60"
Date changed: 11/14/01
Submitted - Wednesday, November 14, 2001 at 19:16:36 (MST)
Upgraded the internet connection to the world to 100mbs. Selected computers (flwo48, linmax) are now at 100mbs. All others now at 10mbs
Ted Groner <ted>
Telescope Involved: all
Date changed: 11/2/2001
Submitted - Wednesday, November 07, 2001 at 16:23:06 (MST)
I installed Black-Box fiber optic isolation to the Hexapod/FLWO48 serial link. I showed Perry how to restore it to origional in the event of failure or flakeyness.
wayne <wayne>
Telescope Involved: 48"
Date changed: 11-1-01
Submitted - Thursday, November 01, 2001 at 14:51:46 (MST)
Instituted Changelog stytem on both 48" and 60"
Ted Groner <ted>
Telescope Involved: all ridge telescopes Date changed: 10/30/01 Submitted- Tuesday, October 30, 2001 at 15:29:29 (MST)